OpenStudy (anonymous):

iff(x)=7x−5andg(x)=2x+3,then g[f(x)]=

OpenStudy (diyadiya):

To find g(f(x)) Replace x in 2x+3 by f(x) that is 7x-5

OpenStudy (diyadiya):

Can you do it?

OpenStudy (anonymous):

Sorry. Can u explain it to me? Because I really dont know how to solve it.

OpenStudy (anonymous):

i dont have an idea about it. Let's say I'm a newbie.

OpenStudy (diyadiya):

You have to find g(f(x)) So you replace x in g(x) by f(x) g(x)=2x+3 f(x)=7x-5 g(f(x))= 2(7x-5) +3 understood till here?

OpenStudy (anonymous):

Owwwww! I know now :) Thankyousomuch!

OpenStudy (diyadiya):

Anytime :)